OpenStudy (anonymous):

what is y=-2/3x+4 in standard form 2x+3y=12 2x+3y=4 3y=-2x+4 3y=-2x+12

OpenStudy (anonymous):

It would be " 2x+3y=12

OpenStudy (anonymous):

All you had to do is run through the problem like this 2x+3y=12 -2x -2x 3y=-2x+12 /3 /3 /3 y=-2/3x+4

OpenStudy (anonymous):

thank you. that helped a lot